# Topological data analysis of the deconfinement transition in SU(3) lattice gauge theory

### v1

@@ Abstract
We study the confining and deconfining phases of pure $\mathrm{SU}(3)$ lattice gauge theory with topological data analysis. This provides unique insights into long range correlations of field configurations across the confinement-deconfinement transition. Specifically, we analyze non-trivial structures in electric and magnetic field energy densities as well as Polyakov loop traces and a Polyakov loop-based variant of the topological density. The Betti curves for filtrations based on the electric and magnetic field energy densities reveal signals of electromagnetic dualities. These dualities can be associated with an interchange in the roles of local lumps of electric and magnetic energy densities around the phase transition. Moreover, we show that plaquette susceptibilities can manifest in the geometric features captured by the Betti curves. We also confront the results obtained for $\mathrm{SU}(3)$ with that for $\mathrm{SU}(2)$ and elaborate on the significant differences. Our results demonstrate that topological data analysis can discriminate phase transitions of different order for non-Abelian lattice gauge theories. Moreover, it provides unprecedented insights into the relevant structures in their vicinity.
@@ I Introduction
Understanding the dynamical mechanism responsible for confinement in non-Abelian gauge theories remains an outstanding challenge. Crucially, the deconfinement phase transition is of second order for only a few simple gauge groups while being first order in the majority of cases, at least in four space-time dimensions. A unifying property among the various proposed confinement mechanisms is the occurrence of topological configurations or defects. These configurations are typically drowned in short range fluctuations, and only become visible via cooling. The latter, however, changes the underlying physics and complicates the access to the underlying dynamics.
@@ III.3 Signals of electromagnetic dualities in Betti curve maxima
The $\beta$-derivative of the expectation value of 11 is given by the plaquette susceptilibility as defined in Eq. (3):
@@ V Conclusions
The maximal Betti numbers also included spatio-temporal geometric manifestations of plaquette trace correlations, which come with a finite-volume peak significantly below the phase transition. Such behavior is in particular possible for first-order phase transitions, less likely for second-order transitions, and thus demonstrates that persistent homology can well distinguish among phase transitions of different order.
@@ Acknowledgements.
Figure 7: Maxima of the Betti curves versus $\beta$ for (a) the $\mathrm{Tr}\,{\mathbf{E}^{2}}$ filtration and (b) the $\mathrm{Tr}\,{\mathbf{B}^{2}}$ filtration and for homology dimensions zero to three from left to right. The displayed flow steps are given by $n_{\mathrm{flow}}\in\{0,1,2,5,10,20\}$; shown colors interpolate between those printed in the legend. Error bars have been computed via jackknife, see Appendix B.
In this appendix, we discuss the influence of cooling on the maximal Betti numbers for the $\mathrm{Tr}\,\mathbf{E}^{2}$ and $\mathrm{Tr}\,\mathbf{B}^{2}$ filtrations displayed in the main text, see Figure 4(a) and (b). For this, in Figure 7 we display the maximal Betti numbers for a range of flow steps $n_{\mathrm{flow}}$ for cooling. We notice that it has a strong influence on both (a) the $\mathrm{Tr}\,\mathbf{E}^{2}$ and (b) the $\mathrm{Tr}\,\mathbf{B}^{2}$ filtrations, whose maximal Betti numbers decrease in values for growing $n_{\mathrm{flow}}$. This is natural for cooling: small-scale fluctuations are increasingly smoothed out with longer cooling times, so the overall number of features is expected to decrease.

### v2

@@ Abstract
We study the confining and deconfining phases of pure $\mathrm{SU}(3)$ lattice gauge theory with topological data analysis. This provides unique insights into long range correlations of field configurations across the confinement-deconfinement transition. Specifically, we analyze non-trivial structures in electric and magnetic field energy densities as well as Polyakov loop traces and a Polyakov loop-based variant of the topological density. The Betti curves for filtrations based on the electric and magnetic field energy densities reveal signals of electromagnetic dualities. These dualities can be associated with an interchange in the roles of local lumps of electric and magnetic energy densities around the phase transition. Moreover, we show that plaquette susceptibilities can manifest in the geometric features captured by the Betti curves. We also compare these findings against earlier results for $\mathrm{SU}(2)$ and elaborate on the significant differences. Our results demonstrate that topological data analysis can identify clear differences between phase transitions of first and second order for non-Abelian lattice gauge theories and provides unprecedented insights into the relevant structures in their vicinity.
@@ I Introduction
Understanding the dynamical mechanism responsible for confinement in non-Abelian gauge theories remains an outstanding challenge. Crucially, the deconfinement phase transition is of second order for only a few simple gauge groups while being first order in the majority of cases, at least in four space-time dimensions Holland et al. (2004, 2003); Pepe and Wiese (2007). A unifying property among the various proposed confinement mechanisms is the occurrence of topological configurations or defects. These configurations are typically drowned in short range fluctuations, and only become visible via cooling. The latter, however, changes the underlying physics and complicates the access to the underlying dynamics.
@@ III.3 Signals of electromagnetic dualities in Betti curve maxima
The $\beta$-derivative of the expectation value of 11 is given by the plaquette susceptibility as defined in Eq. (3):
@@ V Conclusions
The maximal Betti numbers also included spatio-temporal geometric manifestations of plaquette trace correlations, which come with a finite-volume peak significantly below the phase transition. Such behavior is in particular possible for first-order phase transitions but less likely for second-order transitions, and thus demonstrates that persistent homology can identify clear differences between phase transitions of first and second order.
@@ Acknowledgements.
Figure 7: Maxima of the Betti curves versus $\beta$ for homology dimensions zero to three from left to right, where results for the $\mathrm{Tr}\,{\mathbf{E}^{2}}$ filtration are shown in blue and for the $\mathrm{Tr}\,{\mathbf{B}^{2}}$ filtration in orange. Rows (a) through (d) have been computed from samples with different numbers of flow steps $n_{\mathrm{flow}}=1,2,5,10$, respectively. Error bars have been computed via jackknife, see Appendix B. This figure demonstrates the stability of the crossings among the $\mathrm{Tr}\,{\mathbf{E}^{2}}$ and $\mathrm{Tr}\,{\mathbf{B}^{2}}$ filtration Betti curve maxima against cooling.
In this appendix, we discuss the influence of cooling on the maximal Betti numbers for the $\mathrm{Tr}\,\mathbf{E}^{2}$ and $\mathrm{Tr}\,\mathbf{B}^{2}$ filtrations displayed in the main text, see Figure 4(a) to (d). For this, in Figure 7 we display the maximal Betti numbers for a range of flow steps $n_{\mathrm{flow}}$ for cooling. We notice that cooling has a strong influence on both the $\mathrm{Tr}\,\mathbf{E}^{2}$ and the $\mathrm{Tr}\,\mathbf{B}^{2}$ filtrations, whose maximal Betti numbers decrease in values for growing $n_{\mathrm{flow}}$. This is natural for cooling: small-scale fluctuations are increasingly smoothed out with longer cooling times, so the overall number of features is expected to decrease. Yet, the crossings among the maxima Betti numbers for the $\mathrm{Tr}\,{\mathbf{E}^{2}}$ and $\mathrm{Tr}\,{\mathbf{B}^{2}}$ filtrations remain stable against cooling, in particular for the top homology dimension.